# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

### v1

@@ 5 Conclusion & Future Work
In this paper, we introduce a comprehensive benchmark for tabular missing data, MissBench, and a pre-trained transformer for the tabular missing data problem, TabImpute. We build on recent work in tabular representation learning by adapting TabPFN’s architecture and training pipeline for the missing data setting. While we train purely on synthetic data, we are able to impute entries accurately on real-world OpenML data for a comprehensive set of missingness patterns, showcasing our model’s ability to generalize to unseen domains. We open-source our model architecture, weights, and our training and evaluation code (available at https://github.com/jacobf18/tabular). We hope this will facilitate others to utilize, validate, and build on our work.
@@
Also note that since TabImpute uses the same architecture as TabPFN, any improvements to TabPFN’s architecture can be immediately ported to TabImpute. For example, Zeng et al. (2025) and Qu et al. (2025) propose different attention mechanisms to speed up TabPFN-like architectures for tabular data. Finally, since we use a PFN architecture, we output a distribution for each missing entry and can sample from this distribution for multiple imputation (Rubin, 2018). In future work, we plan on (i) exploring further training on more complex missingness patterns and data-generating processes, (ii) enhancing our method to support categorical data, (iii) extending our evaluation set to causal inference settings, which can be modeled as missing-data problems (Agarwal et al., 2023), (iv) improving the architecture to scale to larger datasets, and (v) utilizing our method for multiple imputation.
@@ A.1 Additional tests
Next, we discuss when TabImpute+ does well and when HyperImpute and other methods do well. We found that an important factor in determining performance was the level of missingness. The probability of missingness can only be controlled precisely for MCAR. Thus, we show in Fig. 4 the performance of the top methods as we increase the missingness level.
Figure 4: Imputation accuracy versus fraction of missingness for MCAR. TabImpute+ performs the best when missingness is higher because it is a generative model that fits to the data in context.
@@ A.2.1 Details for NN-MNAR
The pattern simulates a scenario where the propensity $p_{ij}$ depends on the underlying matrix values $X^{*}$ in an arbitrary manner. We achieve a comprehensive coverage of MNAR patterns by leveraging the expressiveness of neural networks.
@@ A.2.2 Details on Seq-MNAR
This pattern simulates a scenario where masking matrix values $M_{ij}\in\{0,1\}$ for each column $j$ are adaptively chosen depending on the information up to column j1j-1 (i.e., regard columns as time). Specifically, we employ variants of bandit algorithms Lattimore & Szepesvári (2020) while regarding the binary masking matrix values as the two arms. Such patterns commonly arise in sequential experiments Ghosh et al. (2024).
@@
The true matrix $X^{*}$ is transformed to constitute the reward. For each designated column $j$, one of the following bandit algorithm utilizes the all the information of $X^{*}$ and $M$ up to column j1j-1 and chooses one of the two arms $\{0,1\}$ via one of the following algorithms: $\varepsilon$-greedy, upper-confidence bound (UCB), Thompson-sampling (Thompson, 1933) or gradient bandit.
@@
We generate exogenous Gaussian noise and add it to the true matrix $X^{*}$ and regard $X^{*}$ as the reward for arm 0 and its noisy version as the reward for arm $1$. Then, starting from the first column with multiple rows as multiple agents, we randomly initiate (with random configurations) one of the four algorithms Lattimore & Szepesvári (2020): $\varepsilon$-greedy, Upper Confidence Bound (UCB) (Auer et al., 2002), Thompson sampling with random configurations (Thompson, 1933). Further, we have the option to randomly mix pooling techniques Ghosh et al. (2024) on top of any of the four algorithms.
@@ A.2.3 Self-Masking-MNAR
This pattern simulates a scenario where the probability of a value being missing is a direct function of the value itself. This pattern can be regarded as a special case of NN-MNAR. A common real-world example includes individuals with very high incomes being less likely to report their salary on a survey.
@@ A.2.4 Censoring-MNAR
This pattern models missingness that arises from the practical/physical limits of measurement equipment, where values below a lower detection limit (left-censoring) or above an upper detection limit (right-censoring) are not recorded. We would expect to see such missingness in sensor or biological assay datasets where test equipment can only detect biomarker levels above/below certain thresholds.
@@ A.2.5 Panel-MNAR
This pattern simulates participant dropout in longitudinal or panel data studies, where once a subject (row) drops out at a specific time point, all their subsequent data is missing. This is a common occurrence in clinical trials or long-term studies/surveys.
@@
For each row $i$, the dropout time $t_{0,i}$ is sampled uniformly from the range of possible time steps, i.e., $t_{0,i}\sim\mathrm{Unif}\{1,...,T-1\}$.
@@ A.2.6 Polarization-MNAR
Values falling in the middle of a feature’s distribution are preferentially removed, simulating survey non-response from individuals with moderate opinions. This is implemented by setting values between the $q$-th and (1q)(1-q)-th quantiles to missing. A ”soft” version makes the observation probability proportional to the value’s distance from the median. Such patterns are most commonly seen in political polls, where moderate voters are less likely to respond while those with extreme views are more likely to respond, or in product reviews, where only very satisfied or very dissatisfied customers leave ratings.
@@ A.2.7 Latent-Factor-MNAR
This pattern generates a complex missingness structure where the probability of an entry being missing depends on unobserved (latent) characteristics of both its row and its column. This is common in recommender systems, where a user’s decision to rate an item depends on latent user preferences and item attributes. Other real-world examples include job application screenings where certain combinations of demographic factors influence whether candidates complete applications, independent of their qualifications, or online dating profiles where cultural background affects disclosure of personal information.
@@ A.2.8 Cluster-MNAR
This pattern induces missingness based on latent group-level characteristics. Rows and columns are first assigned to discrete clusters, and each cluster has a random effect that uniformly influences the observation probability of all its members. This is useful for modeling data from grouped experiments, such as a clinical trial where patients (rows) from a specific hospital (a given row cluster) and certain lab tests (a column cluster) might have systematically different rates of missingness due to local care protocols or equipment availability.
@@ A.2.9 Two-Phase-MNAR
This mechanism mimics multi-stage data collection where a subset of participants from an initial survey (with “cheap” features) are selected for a more detailed follow-up survey (with “expensive” features). The selection for the second phase is dependent on the data collected in the first, making this an MNAR pattern. A specific example includes market research where basic demographics are collected from all participants, but detailed purchasing behavior is only gathered from a subset, with missingness related to income level.
@@ A.3 Additional tables
Column-wise mean (Hawthorne & Elliott, 2005) Mean of columns
@@
SoftImpute (Hastie et al., 2015) Iterative soft thresholding singular value decomposition based on a low-rank assumption on the data
$k$-Nearest Neighbors (Fix & Hodges, 1989) Row-wise nearest neighbors mean
@@
HyperImpute (Jarrett et al., 2022) Iterative imputation method optimizing over a suite of imputation methods
@@
Optimal transport method (Muzellec et al., 2020) Uses optimal transport distances as a loss to impute missing values based on the principle that two randomly drawn batches from the same dataset should share similar data distributions
@@
MissForest (Stekhoven & Bühlmann, 2011) Repeatedly trains a random forest model for each variable on the observed values to predict and fill in missing entries until convergence
@@
ICE (van Buuren & Groothuis-Oudshoorn, 2011) Imputation with iterative and chained equations of linear/logistic models for conditional expectations
@@
MICE (Royston & White, 2011) Handles missing data by iteratively imputing each incomplete variable using regression models that condition on all other variables
@@
GAIN (Yoon et al., 2018) Adapts generative adversarial networks (Goodfellow et al., 2020) where the generator imputes missing values and the discriminator identifies which components are observed versus imputed
@@
MIWAE (Mattei & Frellsen, 2019) Learns a deep latent variable model and then performs importance sampling for imputation
@@
ForestDiffusion (Jolicoeur-Martineau et al., 2024) Trains a diffusion model using XGBoost directly on incomplete tabular data and then fills in missing values with an adapted inpainting algorithm

### v3

@@ 5 Conclusion & Future Work
In this paper, we present a state-of-the-art pre-trained transformer for the tabular missing data problem. We build on recent work in tabular representation learning by adapting TabPFN’s architecture and training pipeline for the missing data setting. Even though we train purely on synthetic data, we are able to impute entries accurately on real-world OpenML data for a comprehensive set of missingness patterns, thus showcasing our model’s ability to generalize to unseen domains. We open-source not only our model architecture and weights, but also our training and evaluation code (available at https://anonymous.4open.science/r/tabular-6F65/README.md). We hope this will facilitate others to validate, utilize, and build upon our work.
@@
Note that since we utilize a base architecture similar to TabPFN, any further improvements to TabPFN’s architecture can be immediately ported to TabImpute. For example, zeng2025tabflex and qu2025tabicl propose different attention mechanisms to speed up TabPFN-like architectures for tabular data. Finally, since we use a PFN architecture, we output a distribution for each missing entry and can sample from this distribution for multiple imputation (rubin2018multiple). In future work, we plan on (i) exploring training further on more complex missingness patterns and data-generating processes, (ii) enhancing our method to impute categorical data better, (iii) extending our evaluation set to causal inference settings, which can be modeled as missing-data problems (agarwal2023causal), (iv) improving the architecture to scale to larger datasets, and (v) utilizing TabImpute further for multiple imputation datasets.
Reproducibility Statement.
We open-source our training code, evaluation code, model architecture and weights, and synthetic data-generation pipeline (available at https://anonymous.4open.science/r/tabular-6F65/). While we utilize 8 H200 GPUs to train our model, our model should be able to fit easily on GPUs with far less memory because our model only has around 22 million parameters. All of the datasets we use for evaluation are public and available in OpenML, making it straightforward to reproduce our results from our pre-trained model.
@@ A.1 Additional tests
Next, we discuss when TabImpute does well and when HyperImpute and other methods do well. We found that an important factor in determining performance was the level of missingness. The probability of missingness can only be controlled precisely for MCAR. Thus, we show in Fig. 4 the performance of the top methods as we increase the missingness level.
Figure 4: Imputation Accuracy vs. probability of missingness for MCAR. TabImpute performs the best when missingness is higher because it is a generative model that fits to the data in-context.
Support for categorical variables.
We support imputing categorical variables via one-hot encodings: First, we convert each categorical column into several one-hot encoding columns. Then, we impute missing entries within this now purely numerical matrix. We then choose the class with the highest score as the categorical imputation for that missing entry. Note that we could also perform a softmax operation over the imputed scores within the one-hot encoded columns to output probabilities over classes. With this method, TabImpute achieves AUC close to that of MissForest and HyperImpute when tested on MCAR missingness with $p=0.4$, as shown in Tab. 4. We leave it for future work to incorporate categorical features into the training procedure, which would likely improve TabImpute’s performance.
Table 4: Area Under the Curve (AUC) Performance on Categorical Columns with MCAR ($p=0.4$) missingness.
OpenML Dataset HyperImpute MissForest TabImpute Mode
ChonicKidneyDisease 0.602 0.582 0.525 0.500
Dog Breeds Ranked 0.566 0.579 0.539 0.500
HappinessRank 2015 0.486 0.466 0.677 0.500
MY DB 0.480 0.514 0.521 0.500
Online Sales 0.790 0.809 0.869 0.500
Parkinson Dataset 0.665 0.640 0.523 0.500
acute-inflammations 0.761 0.756 0.723 0.500
aids 0.511 0.573 0.509 0.500
analcatdata creditscore 0.678 0.672 0.491 0.500
analcatdata cyyoung8092 0.683 0.649 0.611 0.500
analcatdata cyyoung9302 0.713 0.700 0.629 0.500
analcatdata impeach 0.751 0.749 0.706 0.500
analcatdata ncaa 0.549 0.547 0.500 0.500
analcatdata wildcat 0.749 0.682 0.674 0.500
auto price 0.749 0.802 0.761 0.500
backache 0.529 0.516 0.502 0.500
blogger 0.558 0.573 0.499 0.500
caesarian-section 0.534 0.499 0.518 0.500
cloud 0.436 0.409 0.440 0.500
cm1 req 0.677 0.658 0.602 0.500
cocomo numeric 0.629 0.618 0.655 0.500
conference attendance 0.500 0.512 0.499 0.500
corral 0.553 0.558 0.575 0.500
cpu 0.694 0.675 0.546 0.500
fl2000 0.608 0.468 0.533 0.500
flags 0.584 0.583 0.517 0.500
fruitfly 0.602 0.604 0.593 0.500
grub-damage 0.620 0.584 0.541 0.500
hutsof99 logis 0.569 0.568 0.583 0.500
iris 0.885 0.885 0.827 0.500
kidney 0.599 0.460 0.498 0.500
lowbwt 0.587 0.582 0.547 0.500
lung 0.558 0.483 0.496 0.500
lungcancer GSE31210 0.647 0.527 0.567 0.500
lymph 0.641 0.594 0.547 0.500
molecular-biology promoters 0.508 0.505 0.504 0.500
mux6 0.507 0.481 0.502 0.500
nadeem 0.594 0.521 0.542 0.500
nasa numeric 0.596 0.594 0.561 0.500
postoperative-patient-data 0.468 0.494 0.496 0.500
prnn crabs 0.840 0.687 0.737 0.500
prnn viruses 0.563 0.588 0.551 0.500
qualitative-bankruptcy 0.652 0.694 0.645 0.500
servo 0.528 0.516 0.480 0.500
sleuth case1202 0.544 0.590 0.521 0.500
sleuth case2002 0.609 0.564 0.526 0.500
sleuth ex2015 0.589 0.599 0.733 0.500
sleuth ex2016 0.562 0.588 0.532 0.500
tae 0.544 0.592 0.533 0.500
teachingAssistant 0.533 0.528 0.499 0.500
veteran 0.520 0.520 0.502 0.500
white-clover 0.626 0.598 0.622 0.500
zoo 0.754 0.710 0.764 0.500
Overall 0.609 $\pm$ 0.097 0.593 $\pm$ 0.097 0.577 $\pm$ 0.096 0.500 $\pm$ 0.000
@@ A.2.1 Details for NN-MNAR
This pattern simulates a scenario where the propensity $p_{ij}$ depends on the underlying matrix values $X^{*}$ in an arbitrary manner. We achieve a comprehensive coverage of MNAR patterns by leveraging the expressiveness of neural networks.
@@ A.2.2 Details on Seq-MNAR
This pattern simulates a scenario where masking matrix values $M_{ij}\in\{0,1\}$ for each column $j$ are adaptively chosen depending on the information up to column j1j-1 (i.e., regard columns as time). Specifically, we employ variants of bandit algorithms lattimore2020bandit while regarding the binary masking matrix values as the two arms. Such patterns commonly arise in sequential experiments ghosh2024did.
@@
The true matrix $X^{*}$ is transformed to constitute the reward. For each designated column $j$, one of the following bandit algorithm utilizes the all the information of $X^{*}$ and $M$ up to column j1j-1 and chooses one of the two arms $\{0,1\}$ via one of the following algorithms: $\varepsilon$-greedy, upper-confidence bound (UCB), Thompson-sampling (thompson1933likelihood) or gradient bandit.
@@
We generate exogenous Gaussian noise and add it to the true matrix $X^{*}$ and regard $X^{*}$ as the reward for arm 0 and its noisy version as the reward for arm $1$. Then, starting from the first column with multiple rows as multiple agents, we randomly initiate (with random configurations) one of the four algorithms lattimore2020bandit: $\varepsilon$-greedy, Upper Confidence Bound (UCB) (auer2002finite), Thompson sampling with random configurations (thompson1933likelihood). Further, we have the option to randomly mix pooling techniques ghosh2024did on top of any of the four algorithms.
@@ A.2.3 Self-Masking-MNAR
This pattern simulates a scenario where the probability of a value being missing is a direct function of the value itself. This pattern, commonly referred to as ”self-masked MNAR ”, is widely utilized in the missing-data literature to model unbalanced class problems (mohan2018handling; sportisse2020b). A classic example occurs in income surveys, where individuals at the extremes of the income distribution, both very high and very low earners, are often less likely to disclose their income (moore2000income). Similar patterns appear in other domains, such as substance use reporting or medical contexts where patients with severe symptoms might be less able to complete follow-up assessments (ibrahim2009missing).
Within the structured missingness taxonomy proposed by (jackson2023completecharacterisationstructuredmissingness), this mechanism is classified as MNAR-UP (Unstructured, Probabilistic). It is ”Unstructured” because the missingness of an entry $(i,j)$ is conditionally independent of the missingness status of other entries, and ”Probabilistic” as it is determined by a probability function rather than a deterministic rule.
@@ A.2.4 Censoring-MNAR
This pattern models missingness arising from the physical limitations of measurement instruments, where values falling below a lower Limit of Detection or above an upper Limit of Quantification are not recorded. This results in left-censored or right-censored data, a common obstacle in environmental science, epidemiology, and biomedical research (helsel2011statistics; chen2013bayesian). In fields like metabolomics and proteomics, this type of missingness is explicitly recognized as left-censored MNAR (karpievitch2009statistical; lazar2016accounting), as the probability of being missing is directly determined by the concentration falling below the detection threshold. Failing to appropriately account for this, or using simplistic substitution methods like replacing with zero, can introduce substantial bias. Modeling Censoring-MNAR is therefore essential for evaluating imputation methods on datasets generated by instrumentation with inherent sensitivity constraints.
Censoring-MNAR maps to the MNAR-UD (Unstructured, Deterministic) category in the structured missingness framework. It is ”Unstructured” as the missingness of an entry is independent of other missingness indicators, and ”Deterministic” because any value falling beyond the censoring threshold is missing with certainty.
@@ A.2.5 Panel-MNAR
This pattern simulates participant attrition (dropout) in longitudinal and panel data studies. This creates missingness where the dropout of a subject at a specific time point means all subsequent data for that subject is unobserved. Attrition is an inherent feature of long-term studies, including clinical trials and cohort studies (little2019statistical; van2012flexible). The primary concern is ”attrition bias,” which arises when dropout is systematic—meaning those who leave the study differ significantly from those who remain (hausman1979attrition). This often constitutes an MNAR mechanism, particularly in clinical settings where patients might withdraw due to adverse effects or perceived lack of treatment efficacy, directly linking the dropout to the (unobserved) future outcomes (ibrahim2009missing; little1995modeling).
@@
For each row $i$, the dropout time $t_{0,i}$ is sampled uniformly from the range of possible time steps, i.e., $t_{0,i}\sim\mathrm{Unif}\{1,...,T\}$.
@@ A.2.6 Polarization-MNAR
This pattern simulates data where values falling in the middle of a feature’s distribution are preferentially removed, simulating survey non-response from individuals with moderate opinions. This is implemented by setting values between the $q$-th and (1q)(1-q)-th quantiles to missing. A ”soft” version makes the observation probability proportional to the value’s distance from the median. Polarization-MNAR is recognized in the survey methodology literature as a form of voluntary response bias or self-selection bias (bethlehem2010selection). One obvious example can be found in online product reviews, which frequently exhibit a U-shaped or J-shaped distribution, heavily skewed towards the highest and lowest ratings while the middle ground remains sparse—sometimes referred to as the ”brag-and-moan” bias (hu2009overcoming). Similarly, in political polling, nonresponse bias can be exacerbated if highly polarized individuals are more motivated to participate than moderates, potentially exaggerating measures of mass polarization cavari2023survey.
This pattern maps to the ”Unstructured” MNAR categories of (jackson2023completecharacterisationstructuredmissingness). Specifically, our hard polarization variant, where values within a quantile range are missing with certainty, is an instance of MNAR-UD (Unstructured, Deterministic). Our soft polarization variant, where the missingness probability is a function of the value’s distance from the median, is an example of MNAR-UP (Unstructured, Probabilistic).
@@ A.2.7 Latent-Factor-MNAR
This pattern generates a complex missingness structure where the probability of an entry being missing depends on unobserved (latent) characteristics of both its row and its column. This is most common in recommender systems and collaborative filtering, where the data (e.g., user ratings) are inherently MNAR (marlin2009collaborative; steck2010training). Users do not select items to rate uniformly; rather, their exposure to items and their decisions to provide a rating are strongly influenced by their underlying (latent) preferences-the very values the system aims to predict (schnabel2016recommendations). This creates a selection bias where observed ratings are a biased sample of the full data. A standard approach in the literature assumes a low-rank structure for both the underlying data matrix and the missingness mechanism (propensity score matrix), implying they are governed by a shared set of latent factor (sportisse2020aimputation; jin2022matrix). Recent work in causal matrix completion explicitly models these latent confounders (agarwal2023causal).
@@ A.2.8 Cluster-MNAR
This pattern induces missingness based on latent group-level characteristics. Rows and columns are first assigned to discrete clusters, and each cluster has a random effect that uniformly influences the observation probability of all its members. Such structures are prevalent in education (students clustered in schools), healthcare (patients clustered in hospitals), and cross-national surveys. In these settings, missing data patterns are often not independent across observations but rather clustered (van2011multiple). For example, in multi-center clinical trials or Cluster Randomized Trials, missingness can be heavily influenced by site-specific factors such as resource availability or staff training (diaz2014missing). Ignoring this hierarchical structure during imputation fails to account for the within-cluster correlation and can lead to biased estimates (enders2016multilevel). Because the missingness mechanis is related to these cluster-level effects, which may themselves be latent, the data is MNAR.
In the structured missigness taxonomy, we classify Cluster-MNAR as MNAR-UP (Unstructured, Probabilistic) due to the the probabilistic dependence on unobserved factors, without dependence on other missingness indicators.
@@ A.2.9 Two-Phase-MNAR
This pattern mimics multi-stage data collection, mirroring established methodologies such as ”two-phase sampling” (or ”double sampling”) (neyman1938contribution) and ”Planned Missing Data Designs” (graham2006planned). These designs are frequently employed in large-scale surveys and epidemiological studies to manage costs and participant burden when certain variables are expensive or difficult to measure (rhemtulla2016planned). Another example includes market research where basic demographics are collected from all participants, but detailed purchasing behavior is only gathered from a subset, with missingness related to income level. In these designs, a full sample provides baseline information (”cheap” features), and a subset is strategically selected for follow-up (”expensive” features).
@@ A.3 Additional tables
Column-wise mean (hawthorne2005imputing) Mean of columns
@@
SoftImpute (hastie2015matrix) Iterative soft thresholding singular value decomposition based on a low-rank assumption on the data
$k$-Nearest Neighbors (60c19788-1128-3b5f-9275-2d63cc155832) Row-wise nearest neighbors mean
@@
HyperImpute (jarrett2022hyperimpute) Iterative imputation method optimizing over a suite of imputation methods
@@
Optimal transport method (muzellec2020missing) Uses optimal transport distances as a loss to impute missing values based on the principle that two randomly drawn batches from the same dataset should share similar data distributions
@@
MissForest (missforest2011) Repeatedly trains a random forest model for each variable on the observed values to predict and fill in missing entries until convergence
@@
ICE (ice2011) Imputation with iterative and chained equations of linear/logistic models for conditional expectations
@@
MICE (royston2011multiple) Handles missing data by iteratively imputing each incomplete variable using regression models that condition on all other variables
@@
GAIN (gain2018) Adapts generative adversarial networks (goodfellow2020generative) where the generator imputes missing values and the discriminator identifies which components are observed versus imputed
@@
MIWAE (mattei2019miwae) Learns a deep latent variable model and then performs importance sampling for imputation
@@
ForestDiffusion (jolicoeur2024generating) Trains a diffusion model using XGBoost directly on incomplete tabular data and then fills in missing values with an adapted inpainting algorithm
TabPFN The tabpfn-extensions package includes a part to impute missing entries in a table column-by-column using TabPFN.